OpenStudy (anonymous):

Express -3+3i in polar form and find the absolute value

OpenStudy (denisaboichuk):

3i-3

OpenStudy (witty_roger):

magnitude = absolute value=\[\sqrt{3^{2} +(-3)^{2}}\] which is 3sqrt(2)

OpenStudy (witty_roger):

angle = tan inverse of(Imaginary part/Real part) so = atan[(-3)/(3)]=atan(-1) = 180-atan(1)= 180-45=135 degrees

OpenStudy (irishboy123):

always worth drawing these